Suppose you bought a coupon bond today with a face value of $1000, coupon rate is 10%, and maturity period of 15 years. If the interest rate is 8% next period, compute the price of the bond and rate of return if you want to sell it next period. What is the price of the bond if you would like to sell it 5 periods from now?
Answer #1

Face Value =1000
Coupon =10%*1000 =100
Number of Periods left =14
Price of Bond =PV of coupons+PV of Par Value =100*(((1-(1+8%)^-14)/8%)+1000/(1+8%)^14 =1164.88

Price of bond if it sells 5 periods from now =PV of coupons+PV of Par Value
=100*(((1-(1+8%)^-10)/8%)+1000/(1+8%)^10=1134.20